Assume the spot exchange rate is AUD.7064. The expected inflation rate is 1.9 percent in Australia and 1.6 percent in the U.S. What is the expected exchange rate one year from now if relative purchasing power parity exists?

AUD.7063

AUD.7085

AUD.7110

AUD.7074

AUD.7092
Answer #1

Future rate =Spot Rate*(1+Inflation in Australia)/(1+Interest rate) =0.7064*(1+1.9%)/(1+1.6%) =0.7085

Option b is correct option